OpenStudy (anonymous):

The product of three and a squared number is twice the sum of the number and four.

OpenStudy (calculusxy):

Are you trying to find the answer?

OpenStudy (anonymous):

No. Trying to write an algebraic expression for it

OpenStudy (calculusxy):

Ok I will try

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

3n^2 = 2(n+4)

OpenStudy (anonymous):

Thank you!!

OpenStudy (calculusxy):

\[3 \times x^{2}=2(x+4)\]

OpenStudy (anonymous):

You are welcome...read it from left to right and write as you read - easier for me that way